OpenStudy (anonymous):

x2 + 9x + 18 = 0 complete the square to find roots

OpenStudy (anonymous):

x^2 + 2.x. 9/2 + (9/2)^2 - (9/2)^2 + 18 = 0 (x+ 9/2)^2 - 81/4 + 18 = 0 (x + 9/2) ^2 = 9/4 or, (x+9/2) = 3/2 and x + 9/2 = -3/2 x = -3 and x = -6

OpenStudy (anonymous):

what does the "2.x." mean?

OpenStudy (anonymous):

2 * x I just wrote 2*a*b as 2 * x * 9/2

OpenStudy (anonymous):

okay thanks